# TSPDiffuser: Diffusion Models as Learned Samplers for Traveling Salesperson Path Planning Problems

### v1

@@ Abstract
This paper presents TSPDiffuser, a novel data-driven path planner for traveling salesperson path planning problems (TSPPPs) in environments rich with obstacles. Given a set of destinations within obstacle maps, our objective is to efficiently find the shortest possible collision-free path that visits all the destinations. In TSPDiffuser, we train a diffusion model on a large collection of TSPPP instances and their respective solutions to generate plausible paths for unseen problem instances. The model can then be employed as a learned sampler to construct a roadmap that contains potential solutions with a small number of nodes and edges. This approach enables efficient and accurate estimation of traveling costs between destinations, effectively addressing the primary computational challenge in solving TSPPPs. Experimental evaluations with diverse synthetic and real-world indoor/outdoor environments demonstrate the effectiveness of TSPDiffuser over existing methods in terms of the trade-off between solution quality and computational time requirements.
Keywords: Path Planning, Diffusion Model, Traveling Salesperson Problem
1 Introduction
@@
Consider an automated delivery robot tasked with delivering packages to several desks within a cluttered office. To ensure high levels of efficiency and safety, the robot must accomplish two tasks: determine the order of the desks to visit and find collision-free paths between consecutive destinations in obstacle-rich environments. Similar applications may include robotic security patrols [1] and inventory management [2], which we envision will be realized in the near future in unconstrained street corners and stores.
@@
We are interested in a composite problem of traveling salesperson problems (TSPs) [3] and path planning introduced above, which we refer to as traveling salesperson path planning problems (TSPPPs). To solve a TSP for the shortest possible loop that visits all destinations, traveling costs must be estimated in advance for all pairs of destinations via path planning. While various implementations of fast TSP solvers are readily available (e.g., Google OR-Tools Routing Library [4]), path planning algorithms require longer computation times for higher solution quality (i.e., success rate and path optimality). This poses the main computational challenge in solving TSPPPs.
@@
Machine learning has attracted increasing attention as a promising tool for enabling efficient, high-quality path planning in a data-driven manner. Given a large collection of path planning problem instances and their solutions, data-driven path planners learn a neural network to directly generate solution paths [5, 6, 7, 8, 9, 10, 11] or roadmaps/trees on which solution paths can be found efficiently via sampling-based path planning [12, 13, 14, 15, 16]. However, most existing data-driven planners are specialized for a single-query setup (i.e., require the exploration of the same environment for every pair of starts and goals), making them an inefficient choice for TSPPPs that need to execute path planning as many times as the square of the number of destinations.
@@
In this work, we propose TSPDiffuser, a novel data-driven path planner designed to solve TSPPPs. As depicted in Fig. 1, we learn a diffusion model [17, 18] to generate plausible loop paths from obstacle and destination maps. Rather than attempting to generate a single solution path directly from a learned model, TSPDiffuser samples multiple paths to construct a roadmap for a higher success rate and path optimality. Because this roadmap highly likely contains potential solutions with a small number of nodes and edges, it alone can be used to efficiently estimate traveling costs for all destination pairs. In this way, TSPDiffuser aims to achieve both high solution quality and computational efficiency, addressing the main challenge for TSPPPs.
@@
We extensively evaluate our approach using three synthetic environments with diverse obstacle layouts, as well as 30 outdoor and 15 indoor environments with real-world complexity [19, 20, 21]. Experimental results demonstrate that TSPDiffuser significantly improves the trade-off between solution quality and computation time over existing methods [10, 14, 22].
@@
Figure 1: Solving TSPPPs with TSPDiffuser. (a) Problem instance. (b) Generating multiple paths (each shown in different colors) using a learned diffusion model. (c) Connecting paths to construct a roadmap. (d) Perform path planning and solving TSP on the roadmap to obtain a solution path.
2 Related Work
@@
Path planning is a classical problem in AI and robotics, where the goal is to find the lowest-cost feasible path between two states (e.g., positions, poses). Particularly in continuous or high-dimensional state spaces, sampling-based approaches such as probabilistic roadmaps (PRM) [23], rapidly-exploring random trees (RRT) [24], and their asymptotically optimal versions [22] are often used. Recent work aims to leverage machine learning techniques to develop data-driven planners in various ways: learning neural networks that can directly generate paths [5, 6], learning sampling distributions for constructing roadmaps [12, 14, 15], or learning policies and/or value functions for tree-based search [13, 16]. Machine learning modules in such planners are often integrated into classical planning frameworks to ensure high success rates, thus inheriting the trade-off between planning quality and efficiency; the more search costs we spend, the more likely we are to obtain a feasible and optimal solution if one exists. Moreover, most of the existing data-driven planners except [14] consider a single-query setup and need to explore the environment for every pair of sources and destinations.
@@
The traveling salesperson problem (TSP) is a combinatorial optimization problem that aims to find the shortest possible route visiting all given destinations exactly once before returning to the origin, given a set of destinations and pairwise traveling costs among them. TSP has a long history in theoretical computer science and operations research and has been extended to robotics tasks such as order picking in warehouses and vehicle routing [3]. Various forms of TSPs have been studied with practical assumptions on agents (e.g., vehicles, unmanned aerial vehicles, or teams of such moving agents), including Dubins TSP [25], stochastic TSP [26], variable-speed TSP [27], trajectory-based TSP [28], and vehicle routing problems [29]. The proposed TSPPP can be viewed as a specific case of metric TSPs, where traveling costs between destinations are given by the length of shortest possible collision-free paths and thus only satisfy the triangle inequality. Studies in this direction have been relatively limited [30, 31], and no prior work has considered its path planning aspect, i.e., how to accurately and efficiently evaluate traveling costs in obstacle-rich environments.
@@
Diffusion models belong to a family of probabilistic generative models, which are capable of generating high-quality data through an iterative denoising process. This is achieved by progressively injecting noise to disrupt input data, and learning the reverse process that gradually reconstructs the original data back from complete noise (see [18] for an introductory survey.) Diffusion models have led to significant advances especially in image generation [32, 33], and have also begun to be used in robot learning for modeling environmental dynamics [10, 11]. A recent work has attempted to solve Euclidean TSPs using diffusion models [34].
3 Preliminaries
3.1 Problem Definition
Fig. 1 provides an overview of our work. An instance for TSPPPs is represented by an environmental map with obstacles and a set of destinations (Fig. 1a). The solution is expected to be a path that visits all the destinations while avoiding obstacles and returning to the starting point (Fig. 1d). The quality of solution paths is measured by feasibility (i.e., if the path is collision-free) and optimality (i.e., how close the path is to the shortest one.) The performance of planners is evaluated by the trade-off between solution quality and computational efficiency, or in practice, the computation time required to derive solutions.
Our primary objective is to improve this trade-off for solving TSPPPs. Particularly in mobile robotics applications, scalability with respect to the number of destinations can inherently be limited due to battery constraints and the variability of environments. Instead, quick and high-quality path planning becomes more critical to increase the overall throughput in task execution, as it is used thoroughly to estimate traveling costs between destinations.
@@ 3.1 Problem Definition
3.2 Denoising Diffusion Probabilistic Models (DDPMs)
@@ 4.1 Model Architecture
Specifically, we represent obstacles and destinations by binary images and their distance images obtained via distance transform. These images are stacked and encoded into a $d$-dimensional unified feature map $F$ using fully-convolutional networks. The feature vector extracted at location $p$ from the feature map, denoted as $F(p)\in\mathbb{R}^{d}$, describes a local context about whether there are obstacles and destinations nearby. Our noise estimator takes the form of $\varepsilon_{\theta}(\tau^{i},i,F(\tau^{i}))$, where $F(\tau^{i})=(F(p^{i}_{1}),\dots,F(p^{i}_{T}))$ is a sequence of feature vectors extracted along $\tau^{i}=(p^{i}_{1},\dots,p^{i}_{T})$. In practical implementations, we concatenate the position $p^{i}_{t}$, the denoising timestep $i$ encoded into a vector via Fourier transform, and the feature vector $F(p^{i}_{t})$ to form the input to the noise estimation network.
4.2 Constructing Roadmaps using Learned Diffusion Models
Despite their high generation capability, we observe two limitations for diffusion models when used in path planning. Firstly, sample generation by diffusion models is typically computationally intensive, which is not necessarily suitable for practical applications where higher throughput with limited computation resources is appreciated. Although many extensions have been developed to mitigate this problem (e.g., [36, 37]), efficient sampling often sacrifices the quality of generated samples. Secondly, even with reward-guided sampling, there is no guarantee that generated paths satisfy constraints imposed in TSPPPs, i.e., visiting all destinations in a single loop, while avoiding collisions with any single obstacle. This makes it hard to solve TSPPPs by diffusion models alone.
@@ 4.2 Constructing Roadmaps using Learned Diffusion Models
Alg. 1 describes the proposed algorithm. Let $\{\tau_{1},\dots,\tau_{M}\}$ be a set of paths generated from a learned diffusion model, where $\tau_{m}=(p_{m,1},\dots,p_{m,T_{m}})$ is a sequence of positions. All the positions contained in the generated paths are regarded as nodes $V$ (L3). With a validity check function $\texttt{isvalid}(\cdot,\cdot;\Omega_{\mathrm{free}})$ that returns true if there are no obstacles between a pair of positions, we first connect nodes along each generated path to define edges $E$ (L4-L7). After adding destinations $\Omega_{\mathrm{dst}}$ and nodes around obstacle boundaries $\{b_{1},\dots,b_{N}\}$ to $V$ (L9), we further span edges between each node and its $K$ nearest neighbors to $E$ (L10-L15).
@@
11:$V\leftarrow V\cup\Omega_{\mathrm{dst}}\cup\{b_{1},\dots,b_{N}\}$triangleright Add destinations and $N$ boundary nodes
@@
4.3 Solving TSPPP
On the constructed roadmap, we perform path planning to calculate the shortest-path distance as traveling costs between each pair of destinations. We then solve a TSP using the calculated traveling costs. This gives us the order of destinations to visit with the total path cost minimized among all possible tours. Finally, we stitch the shortest paths between destinations in the determined order to generate the solution path for TSPPPs.
5 Experiments
5.1 Planning Environments
We extensively evaluate TSPDiffuser using both synthetic and real-world indoor/outdoor navigation environments, each with various layouts of obstacles and destinations on 128x128 sized maps:
  • Synthetic: Three different synthetic environments with different obstacle layouts; Standard with twenty random 20x20 square-shaped obstacles, More Obstacles with forty random 10x10 square-shaped obstacles, and Larger Obstacles with ten random 40x40 square-shaped obstacles.
@@ 5.1 Planning Environments
For each environment, we created 100 instances by randomizing the locations of obstacles (for Standard, More Obstacle, and Larger Obstacle) and cropped areas (for City/Street Map and Indoor Navigation Dataset). For each instance, ten random destinations were sampled from the largest non-obstacle area. The ground-truth, i.e., (pseudo-) optimal, path for each instance was obtained by computing traveling costs among destinations using PRM [22] with 3,000 nodes and solving the TSP using or-tools [4].
In addition, we generated approximately 100,000 instances of the Standard environment as a training dataset for TSPDiffuser and other data-driven baselines presented later. This verifies the generalization ability of these planners, especially from synthetic to real-world environments. Ground-truth paths in the training dataset were repeated until the total number of temporal steps exceeded 256, and then trimmed to have a fixed step of 256. This process allowed us to easily stack multiple ground-truth paths to form a mini-batch during model training.
5.2 Evaluation Setups
@@
For each environment, we calculated the success rate weighted by path length (SPL) [38] $\frac{1}{100}\sum_{i=1}^{100}S_{i}\frac{\hat{l}_{i}}{\max(l_{i},\hat{l}_{i})}$ for 100 instances, where $S_{i}$ is the binary indicator that takes $S_{i}=1$ if the method found a solution path for the $i$-th instance and $S_{i}=0$ otherwise, and $l_{i},\hat{l}_{i}$ are the lengths of the solution and ground-truth paths. We evaluate the trend of SPL scores while changing hyper-parameters of a method that affect the trade-off between the solution quality and required computation times, such as the number of nodes for PRMs and that of generated paths $M$ for TSPDiffuser. All the evaluation procedures were run on the Intel® Xeon® CPU @ 2.20GHz.
@@
Due to the absence of prior work that solves the same problem, we have extended existing path planning algorithms for baseline methods. All the baselines below are able to construct a roadmap to evaluate traveling costs for all combinations of destinations (i.e., support multi-query setups), which is then followed by the same solution procedure shown in Sec. 4.3.
@@
  • Diffuser [10]: A state-of-the-art trajectory generation method using diffusion models. For a fair comparison, we employed the same diffusion model architecture as that of TSPDiffuser in Sec. 4.1. A single loop path is generated from a learned model and then connected with destinations and boundary nodes to construct a roadmap. This baseline can also be viewed as a degraded version of TSPDiffuser where the number of paths is limited to $M=1$.
@@
TSPDiffuser consists of the problem instance encoder network and the noise estimation network. As the instance encoder, we used a U-Net [35] with four encoding and decoding layers from the VGG backbone [39], which outputs $d=256$-dimensional feature maps. The noise estimation network was a 1D U-Net with residual blocks and self-attention modules, where the input is the sequence of concatenations of positions, denoising timesteps encoded into 16-dimensional real-valued vectors via Fourier transform, and the 256-dimensional feature vectors extracted from the feature maps. These two modules were trained end-to-end with the standard DDPM training pipeline [17] for 300 epochs with a batch size of 32, using the AdamW optimizer [40] with the cosine annealing learning rate scheduler [41]. The granularity of denoising timesteps, i.e., $I$ in Sec. 3.2, was set to 500 for training, and 5 for sampling. For roadmap construction, the number of sampled paths $M$ were set to either one of $M\in\{5,10,20,30\}$, where larger $M$ will result in higher SPL scores with more computation times. The reward guidance, i.e., $\nabla\mathcal{J}(\tau)$ in Eq. (3), is given by the sum of distance images from obstacles and destinations, which guides output positions to avoid obstacles and approach destinations. The number of neighbors considered when spanning edges among paths, i.e., $K$ in Alg. 1, were set to $K=5$ unless specified otherwise. The number of boundary nodes, i.e., $N$ in Alg. 1, were set adaptively by detecting boundary pixels of obstacles.
5.3 Results

### v2

@@ Abstract
This paper presents TSPDiffuser, a novel data-driven path planner for traveling salesperson path planning problems (TSPPPs) in environments rich with obstacles. Given a set of destinations within obstacle maps, our objective is to efficiently find the shortest possible collision-free path that visits all the destinations. In TSPDiffuser, we train a diffusion model on a large collection of TSPPP instances and their respective solutions to generate plausible paths for unseen problem instances. The model can then be employed as a learned sampler to construct a roadmap that contains potential solutions with a small number of nodes and edges. This approach enables efficient and accurate estimation of travel costs between destinations, effectively addressing the primary computational challenge in solving TSPPPs. Experimental evaluations with diverse synthetic and real-world indoor/outdoor environments demonstrate the effectiveness of TSPDiffuser over existing methods in terms of the trade-off between solution quality and computational time requirements.
I Introduction
@@
Imagine an automated delivery robot tasked with delivering packages to several desks within a cluttered office. To ensure high levels of efficiency and safety, the robot must accomplish two tasks: determine the order of the desks to visit and find collision-free paths between consecutive destinations in obstacle-rich environments. Similar applications may include robotic security patrols [1] and inventory management [2], which we envision will be realized in the near future in unconstrained street corners and stores.
@@
We are interested in a composite problem of traveling salesperson problems (TSPs) [3] and path planning introduced above, which we refer to as traveling salesperson path planning problems (TSPPPs). To solve a TSP for the shortest possible loop that visits all destinations, travel costs must be estimated in advance for all pairs of destinations via path planning. While various implementations of fast TSP solvers are readily available (e.g., Google OR-Tools Routing Library [4]), path planning algorithms require longer computation times for higher solution quality (i.e., success rate and path optimality). This poses the main computational challenge in solving TSPPPs, especially when the obstacle layout in the environment can change regularly.
@@
Machine learning has attracted increasing attention as a promising tool for enabling efficient, high-quality path planning in a data-driven manner. Given a collection of path planning problem instances and their solutions, data-driven path planners learn a neural network to directly generate solution paths [5, 6, 7, 8, 9, 10] or roadmaps/trees on which solution paths can be found efficiently via sampling-based path planning [11, 12, 13, 14, 15]. However, most existing data-driven planners are specialized for a single-query setup, i.e., require the exploration of the same environment for every pair of starts and goals. This makes these planners an inefficient choice for TSPPPs that need to execute path planning as many times as the square of the number of destinations for every obstacle layout.
@@
In this work, we propose TSPDiffuser, a novel data-driven path planner designed to solve TSPPPs. At its core, we learn a diffusion model [16, 17] to generate plausible loop paths from obstacle and destination maps. Rather than attempting to generate a single solution path directly from a learned model, TSPDiffuser samples multiple paths to construct a roadmap for a higher success rate and path optimality as depicted in Fig. 1. Because this roadmap highly likely contains potential solutions with a small number of nodes and edges, it alone can be used to efficiently estimate travel costs for all destination pairs. In this way, TSPDiffuser aims to achieve both high solution quality and computational efficiency, addressing the main challenge for TSPPPs.
@@
We extensively evaluate our approach using three synthetic environments with diverse obstacle layouts, as well as 30 outdoor and 15 indoor environments with more complex layouts of obstacles scanned from real-world buildings, walls, or furniture [18, 19, 20]. Experimental results demonstrate that TSPDiffuser significantly improves the trade-off between solution quality and computation time over existing methods [9, 13, 21].
Figure 1: Solving TSPPPs with TSPDiffuser (right). We leverage a diffusion model as a learned sampler to construct a roadmap for traveling salesperson path planning problems (TSPPPs). Our approach significantly improves the balance between solution quality and efficiency compared to existing sampling-based path planners such as probabilistic roadmaps (PRM; left)
@@
Figure 2: Solution Overview. (a) Problem instance. (b) Generating multiple paths (each shown in different colors) using a learned diffusion model. (c) Connecting paths to construct a roadmap. (d) Perform path planning and solving TSP on the roadmap to obtain a solution path.
II Related Work
@@
Path planning is a classical problem in AI and robotics, where the goal is to find the lowest-cost feasible path between two states (e.g., positions, poses). Particularly in continuous or high-dimensional state spaces, sampling-based approaches such as probabilistic roadmaps (PRM) [22], rapidly-exploring random trees (RRT) [23], and their asymptotically optimal versions [21] are often used. Recent work aims to leverage machine learning techniques to develop data-driven planners in various ways: learning neural networks that can directly generate paths [5], learning sampling distributions for constructing roadmaps [11, 13, 14], or learning policies and/or value functions for tree-based search [12, 15]. Machine learning modules in such planners are often integrated into classical planning frameworks to ensure high success rates, thus inheriting the trade-off between planning quality and efficiency; the more search costs we spend, the more likely we are to obtain a feasible and optimal solution if one exists. Moreover, most of the existing data-driven planners except [13] consider a single-query setup and need to explore every environment for every pair of sources and destinations.
@@
TSP is a combinatorial optimization problem that aims to find the shortest possible route visiting all given destinations exactly once before returning to the origin, given a set of destinations and pairwise travel costs among them. TSP has a long history in theoretical computer science and operations research and has been extended to robotics tasks such as order picking in warehouses and vehicle routing [3]. Recent work has attempted to extend machine learning techniques to solve Euclidean TSPs [24, 25, 26, 27]. Various other forms of TSPs have also been studied with practical assumptions on agents (e.g., vehicles, unmanned aerial vehicles, or teams of such moving agents), including Dubins TSP [28], stochastic TSP [29], variable-speed TSP [30], trajectory-based TSP [31], and vehicle routing problems [32]. The proposed TSPPP can be viewed as a specific case of metric TSPs, where travel costs between destinations are given by the length of shortest possible collision-free paths and thus only satisfy the triangle inequality. Studies in this direction have been relatively limited [33, 34]. No prior work has considered its path planning aspect, i.e., how to accurately and efficiently evaluate travel costs in obstacle-rich environments [^1] In parallel to this work, [35] has recently studied multi-goal path finding, a problem setting very close to TSPPPs..
@@
Diffusion models belong to a family of probabilistic generative models, which are capable of generating high-quality data through an iterative denoising process. This is achieved by progressively injecting noise to disrupt input data, and learning the reverse process that gradually reconstructs the original data back from complete noise (see [17] for an introductory survey.) Diffusion models have led to significant advances especially in image generation [36, 37], and have also begun to be used in robot learning for modeling environmental dynamics [9, 10].
III Preliminaries
III-A Problem Definition
Fig. 2 provides the overview of our work. An instance for TSPPPs is represented by an environmental map with obstacles and a set of destinations (Fig. 2a). The solution is expected to be a path that visits all the destinations while avoiding obstacles and returning to the starting point (Fig. 2d). The quality of solution paths is measured by feasibility (i.e., if the path is collision-free) and optimality (i.e., how close the path is to the shortest one.) The performance of planners is evaluated by the trade-off between solution quality and computational efficiency, or in practice, the computation time required to derive solutions.
Our primary objective is to improve this trade-off for solving TSPPPs. Particularly in mobile robotics applications, scalability with respect to the number of destinations can inherently be limited due to battery constraints. Instead, quick and high-quality path planning becomes more critical to increasing overall task throughput, because it must be used thoroughly to estimate travel costs between destinations for every change in the obstacle layout in the environment.
@@ III-A Problem Definition
III-B Denoising Diffusion Probabilistic Models (DDPMs)
@@ IV-A Model Architecture
Specifically, we represent obstacles and destinations by binary images and their distance images obtained via distance transform. These images are stacked and encoded into a $d$-dimensional unified feature map $F$ with fully-convolutional networks. The feature vector extracted at location $p$ from the feature map, denoted as $F(p)\in\mathbb{R}^{d}$, describes a local context about whether there are obstacles and destinations nearby. Our noise estimator takes the form of $\varepsilon_{\theta}(\tau^{i},i,F(\tau^{i}))$, where $F(\tau^{i})=(F(p^{i}_{1}),\dots,F(p^{i}_{T}))$ is a sequence of feature vectors extracted along $\tau^{i}=(p^{i}_{1},\dots,p^{i}_{T})$. In practical implementations, we concatenate the position $p^{i}_{t}$, the denoising timestep $i$ encoded into a vector via Fourier transform, and the feature vector $F(p^{i}_{t})$ to form the input to the noise estimator.
IV-B Constructing Roadmaps using Learned Diffusion Models
Despite their high generation capability, we observe two limitations for diffusion models when used in path planning. Firstly, there is an inevitable trade-off between the quality of generated samples and the required computation time. Although higher throughput with limited computational resources is desirable for practical robotic applications, efficient sampling often sacrifices the sample quality. Secondly, even with reward-guided sampling, there is no guarantee that generated paths satisfy constraints imposed in TSPPPs, i.e., visiting all destinations in a single loop, while avoiding collisions with any single obstacle. This makes it hard to solve TSPPPs by diffusion models alone.
@@ IV-B Constructing Roadmaps using Learned Diffusion Models
Alg. 1 describes the proposed algorithm. Let $\{\tau_{1},\dots,\tau_{M}\}$ be a set of paths generated from a learned diffusion model, where $\tau_{m}=(p_{m,1},\dots,p_{m,T_{m}})$ is a sequence of positions. All the positions contained in the generated paths are regarded as nodes $V$ (L3). With a validity check function $\texttt{isvalid}(\cdot,\cdot;\Omega_{\mathrm{free}})$ that returns true if there are no obstacles between a pair of positions, we first connect nodes along each generated path to define edges $E$ (L4-L7). After adding destinations $\Omega_{\mathrm{dst}}$ and nodes around boundaries of the obstacles collided with generated paths, $\{b_{1},\dots,b_{N}\}$, to $V$ (L9), we further span edges between each node and its $K$ nearest neighbors to $E$ (L10-L15).
@@
11:$V\leftarrow V\cup\Omega_{\mathrm{dst}}\cup\{b_{1},\dots,b_{N}\}$triangleright Add destinations and $N$ boundary nodes for the obstacles collided with generated paths
@@
IV-C Solving TSPPP
On the constructed roadmap, we perform path planning to calculate the shortest-path distance as a travel cost between pairs of destinations. We then solve a TSP using the calculated travel costs. This gives us the order of destinations to visit with the total path cost minimized among all possible tours. Finally, we stitch the shortest paths between destinations in the determined order to generate the solution for TSPPPs.
V Experiments
V-A Planning Environments
We extensively evaluate TSPDiffuser using the following synthetic and real-world indoor/outdoor navigation environments, each with various layouts of obstacles and destinations on 128x128 sized maps:
  • Synthetic: Three synthetic environments with different obstacle layouts; Standard with twenty random 20x20 square-shaped obstacles, More Obstacles with forty random 10x10 square-shaped obstacles, and Larger Obstacles with ten random 40x40 square-shaped obstacles.
@@ V-A Planning Environments
Note that obstacles in the City/Street Map and Indoor Navigation Dataset are much more complex in their shapes and layouts and faithful to real-world settings, as they are the scans of real-world buildings, walls, or furniture. For each environment, we created 100 instances by randomizing the locations of obstacles for the Synthetic environments and cropped areas for City/Street Map and Indoor Navigation Dataset. For each instance, ten random destinations were sampled from the largest non-obstacle area. The ground-truth, i.e., (pseudo-) optimal, path for each instance was obtained by computing travel costs among destinations using PRM [21] with 3,000 nodes and solving the TSP using or-tools [4].
In addition, we generated approximately 100,000 instances of the Standard environment as a training dataset for TSPDiffuser and other data-driven baselines presented later. This verifies the generalization ability of these planners from synthetic to real-world environments (i.e., sim-to-real transfer ability). Ground-truth paths in the training dataset were repeated until the total number of temporal steps exceeded 256, and then trimmed to have a fixed step of 256. This process allowed us to easily stack multiple ground-truth paths to form a mini-batch during training.
V-B Evaluation Setups
@@
For each environment, we calculated the success rate weighted by path length (SPL) [39] for 100 instances: $\frac{1}{100}\sum_{i=1}^{100}S_{i}\frac{\hat{l}_{i}}{\max(l_{i},\hat{l}_{i})}$. Here, $S_{i}$ is the binary indicator that takes $S_{i}=1$ if the method found a solution path for the $i$-th instance and $S_{i}=0$ otherwise, and $l_{i},\hat{l}_{i}$ are the lengths of the solution and ground-truth paths. As such, SPL is used to evaluate both success rate and path optimality in a single metric. We evaluate the trend of SPL scores while changing hyper-parameters of a method that affect the trade-off between the solution quality and required computation times, such as the number of nodes for PRMs and that of generated paths $M$ for TSPDiffuser. All the evaluation procedures were run on the Intel® Xeon® CPU @ 2.20GHz.
@@
Due to the absence of prior work that solves the same problem, [^2] Note that existing learning-based TSP solvers [24, 25, 26, 27] leverage machine learning to solve TSPs while assuming that pairwise travel costs are already given, thus cannot be compared to the proposed method. we have extended existing path planning algorithms for baseline methods. All the baselines below are able to construct a roadmap to evaluate travel costs for all combinations of destinations (i.e., support multi-query setups), which is then followed by the same solution procedure shown in Sec. IV-C.
@@
  • Diffuser [9]: A state-of-the-art trajectory generation method using diffusion models. For a fair comparison, we employed the same diffusion model architecture as that of TSPDiffuser in Sec. IV-A. A single loop path is generated from a learned model and then connected with destinations and boundary nodes to construct a roadmap. This baseline can be viewed as a degraded version of TSPDiffuser where the number of paths is $M=1$.
@@
For fairness in evaluation, all methods are implemented on a single codebase written in python, with the same function shared for edge validity check (i.e., $\texttt{isvalid}(\cdot,\cdot;\Omega_{\mathrm{free}})$ in Sec. IV-B.) TSPDiffuser consists of the problem instance encoder network and the noise estimation network. As the instance encoder, we used a U-Net [38] with four encoding and decoding layers from the VGG backbone [40], which outputs $d=256$-dimensional feature maps. The noise estimation network was a 1D U-Net with residual blocks and self-attention modules, where the input is the sequence of concatenations of positions, denoising timesteps encoded into 16-dimensional real-valued vectors via Fourier transform, and the 256-dimensional feature vectors extracted from the feature maps. These two modules were trained end-to-end with the standard DDPM training pipeline [16] for 300 epochs with a batch size of 32, using the AdamW optimizer with the cosine annealing learning rate scheduler. The granularity of denoising timesteps, i.e., $I$ in Sec. III-B, was set to 500 for training, and 5 for sampling. For roadmap construction, the number of sampled paths $M$ were set to either one of $M\in\{5,10,20,30\}$, where larger $M$ will result in higher SPL scores with more computation times. The reward guidance, i.e., $\nabla\mathcal{J}(\tau)$ in Eq. (3), is given by the sum of distance images from obstacles and destinations, which guides output positions to avoid obstacles and approach destinations. The number of neighbors considered when spanning edges among paths, i.e., $K$ in Alg. 1, were set to $K=5$ unless specified otherwise. The number of boundary nodes, i.e., $N$ in Alg. 1, were set adaptively by detecting boundary pixels of obstacles.
V-C Results